OpenStudy (anonymous):

Justify the property: 1. If a/b = c/d, then b/a = d/c 2. If a/b = c/d, then a+b/b = c+d/d Thanks! x

OpenStudy (raden):

1. a/b = c/d cross multiply, get bc = ad divided by ac on both sides, get bc/ac = ad/ac (cancel out the similar terms) b/a = d/c

OpenStudy (raden):

2. If a/b = c/d add by 1 on both sides, get a/b + 1 = c/d + 1 a/b + b/b = c/d + d/d (a+b)/b = (c+d)/d

OpenStudy (anonymous):

Thanks so much! Is there an actual "name" for the properties? (ex: Transitive property)